OpenStudy (anonymous):

Solve using substitution y=2x=3 y=3x+1

OpenStudy (anonymous):

y=2x-3 y=3x+1 (2x-3)-3 (3x+1)2 -6x+9 6x+2 y=11 3(11)+1 33+1 x=34 i think this is right